OpenStudy (mayaal):

giving out medals!!! A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 576t. After how many seconds does the projectile take to reach its maximum height? Show your work for full credit.

OpenStudy (anonymous):

h(t) = -16t^2 + 640t This is a "quadratic" (think parabola) and because the coefficient associated with the t^2 term is NEGATIVE, we know it opens downward. Therefore, the vertex is the maximum. . The time (secs) when it reaches the max is: t = -b/(2a) t = -640/(2(-16)) t = -640/(-32) t = 20 seconds

OpenStudy (anonymous):

this is just an example do the same thing with ur problem

OpenStudy (anonymous):

medal, hello where is it

OpenStudy (mayaal):

thankyou,but are we are looking for the vertex,aren't we?

OpenStudy (anonymous):

yes yes see if this link helps http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.644714.html

OpenStudy (mayaal):

oh I got it.thankyou again:)

OpenStudy (anonymous):

welcome:)